# How (Mis)calibrated is Your Federated CLIP and What To Do About It?

### v1

@@ 1 Introduction
Reliability is a critical consideration when deploying deep learning (DL) models in production scenarios. We aim to develop DL models that are not only highly accurate but also reliable. Reliability has many facets, including robustness to covariate shift [csurka2017domain], model calibration [guo2017calibration], and generalization to out-of-distribution data [yang2024generalized]. Of particular interest, model calibration is crucial because it ensures that a model’s predicted probabilities accurately reflect the true likelihood of correctness, enabling reliable decision-making [pavlovic2025understanding]. As powerful DL models (e.g., CLIP [radford2019language]) are increasingly being deployed in high-stake applications, such as healthcare and autonomous systems, ensuring proper calibration becomes crucial.
Table 1: Overview of works focusing on model calibration. While previous works have studied calibration of CLIP, no work has studied how fine-tuning CLIP in the federated learning (FL) setting affects calibration. Our work explores this direction.
Method CLIP Calibration Fine-tune Offline FL
@@
Figure 1: Comparison of reliability diagrams and calibration errors. Expected Calibration Error (ECE $\downarrow$) [guo2017calibration] is reported on the OxfordPets [oxfordpets] dataset on a centralized training or offline (left) and a non-IID personalized federal learning (FL) setting (right).
@@
In this work, we make an attempt to answer the question: how (mis)calibrated is CLIP when fine-tuned in a FL setup? Towards that goal, we first analyze if the trend of miscalibration, observed in the offline CLIP fine-tuning setting [dor, wang2024open], also holds true in the FL setting. As shown in Fig. 1, CoOp [coop], a textual-prompt tuning approach, when fine-tuned in the FL setup (PromptFL [promptfl]) is even more miscalibrated. Imposing regularization through in-training calibration (e.g., with DCA [dca]) to both the offline and FL setting alleviates miscalibration by some extent. This observation conforms with the findings of Chu et al. [nucfl], who demonstrated a similar behaviour in the FL setup. However, as shown in Tab. 1, the work in [nucfl] only experimented with the classical ResNet family of models trained from scratch and did not explore a multimodal setting. Differently, in our study we are specifically interested in analyzing the miscalibration of fine-tuned CLIP models in the FL setup, which exhibits substantially different dynamics from standard Convnets [minderer2021revisiting]. In Sec. 4, we dive deep into this issue, by investigating other factors for improving calibration, such as incorporating existing calibration techniques, and global aggregation strategies in FL, and find consistent trends.
@@
Although applying extra regularization loss for improving calibration is effective in lowering calibration error, it introduces additional computation and design considerations. Therefore, we strive to find an approach that not only preserves the accuracy of the model, but also ensures a well-calibrated model without introducing additional regularization. To this end, we deviate from the common approach of textual prompt-tuning in FL [promptfl, diprompt, fedotp], and adopt low rank adaptation (LoRA [lora]) as a means to fine-tune CLIP. We call our approach $\text{FL}^{2}\text{oRA}$ (Federated Learning with LoRA). As shown in Fig. 1 (right), $\text{FL}^{2}\text{oRA}$ alone is better calibrated than PromptFL + DCA, without the need for any explicit regularization on the network.
We further hypothesize that the core problem of miscalibration in the FL setting is not merely how we aggregate or calibrate the CLIP base model, but what we fine-tune. LoRA due to its property of introducing minimal variations, compounded with the regularization effect induced by FedAvg [fedavg], results in a strongly calibrated model (see Sec. 4.4 for a detailed discussion). Note that the goal of this work is not to introduce a novel architecture for FL, but to demonstrate that an appropriately chosen parameter-efficient finetuning technique (PEFT), i.e., LoRA, is sufficient to preserve calibration. Our empirical evaluation in Sec. 4 spanning across three different settings under varying levels of data heterogeneity confirms that fine-tuning with LoRA alone is highly effective in preserving model calibration in the FL setting, and adding extra regularization to improve calibration brings only a marginal improvement.
@@
Contributions. In summary: (i) We are the first to study the impact of fine-tuning CLIP on the model’s calibration in the FL setup. Our study reveals that such models, without any extra calibration, are miscalibrated, and thus deserve more attention to improve their calibration. (ii) We present $\text{FL}^{2}\text{oRA}$, a simple and effective approach, that departs from the dominant approach of fine-tuning CLIP with textual prompt-tuning. (iii) We perform extensive experiments across diverse FL scenarios and demonstrate that our LoRA-based federated fine-tuning approach consistently serves as a strong baseline, potentially eliminating the need of additional regularization for improved calibration.
@@ 2 Related Work
Federated Learning (FL) enables collaborative model training across decentralized clients without sharing local data. A key challenge lies in client heterogeneity, e.g., data on different clients is typically non-IID, which negatively affect performance. Previous research on FL primarily focused on developing robust aggregation and optimization strategies to mitigate these issues. Representative approaches include FedAvg [fedavg], the reference FL algorithm that averages client model updates and FedProx [fedprox], which adds a proximal term to each client’s local objective to limit client drift under heterogeneity. Other methods improve over FedAvg with respect to convergence using control variates to reduce variance caused by non-IID updates [scaffold] or modifying local objectives with dynamic regularizers. The main objective of these works is similar: improve the quality of the global model (e.g., accuracy, convergence) obtained by aggregating heterogeneous client updates.
In this work, we study an underexplored aspect of FL, i.e. how federated training affects model calibration in VLMs.
@@
Federated Fine-tuning of CLIP. When applied to large VLMs like CLIP [radford2019language]], FL faces major challenges, as fine-tuning millions of parameters implies high computational and communication overhead during client training and aggregation. To alleviate these problems, prompt tuning [coop, kgcoop, vpt, maple] has emerged as an efficient alternative for adapting CLIP. Several works have extended prompt tuning to the federated setting [bai2024diprompt, promptfl, qiu2024federated, sun2024towards, fedotp, fedtpg, fedpgp, fedpha]. For instance, PromptFL [promptfl] adapts prompt learning to decentralized clients by introducing regularization across client prompts. FedOTP [fedotp] and FedTPG [fedtpg] enhance communication efficiency and personalization through orthogonal or task-conditioned prompt generation. FedCLIP [fedclip] decouples the global and local adaptation of text and image encoders. Methods such as FedTPG [fedtpg] and FedMVP [fedmvp] dynamically generate prompts conditioned on auxiliary information (i.e. class names or visual attributes).
@@
Differently, in this work we propose to adopt a PEFT approach, i.e. LoRA [biderman2024lora], to adapt CLIP in a federated setting. Although FL with LoRA has been explored in the context of Large Language Models [yang2025federated], prior work has focused mainly on accuracy rather than calibration, and to our knowledge, ours is the first to study model calibration in CLIP-based federated image classification.
@@
Model Calibration [pavlovic2025understanding] measures how well a model predicted confidence scores reflect the true likelihood of correctness. Calibration methods are typically categorized into post-hoc approaches and methods that operate at training time. Post-hoc methods, such as temperature scaling [guo2017calibration] and Dirichlet calibration [kull2017beta], adjust the model’s output probabilities after training without altering the model parameters, offering simplicity and efficiency but limited flexibility. The latter category of methods modify the learning process itself to encourage well-calibrated predictions. Examples include label smoothing [muller2019does], which softens target labels, focal loss [lin2017focal], which emphasizes harder-to-classify samples and methods like DCA [dca] and MDCA [hebbalaguppe2022stitch], which adjust the training objective to optimize calibration metrics. These methods often achieve better calibration than post-hoc approaches but may require careful hyperparameter tuning.
@@
In the context of VLMs, zero-shot CLIP is known to be well-calibrated [minderer2021revisiting]. However, fine-tuning can negatively impact calibration, prompting recent solutions like DOR [dor], DAC [wang2024open], and CalShift [khan2025confidence] to restore reliable confidence estimates. Most calibration studies have been conducted in the offline setting. In the FL setting, only a few works [peng2024fedcal, nucfl] have addressed calibration, primarily using convolutional networks that are fully fine-tuned. To the best of our knowledge, no prior work has explored the calibration of CLIP-based FL methods.
@@ 3 Preliminaries
In this section, we discuss the preliminaries related to federated learning (Sec. 3.1), federated fine-tuning of CLIP (Sec. 3.2) and model calibration (Sec. 3.3).
3.1 Federated Learning
For the task of image classification, the FL scheme consists of training data scattered among $N$ clients and each client $n$ possess a dataset $\mathcal{D}_{n}=\{\mathbf{x}^{(i)},y^{(i)}\}^{\mid D_{n}\mid}_{i=1}$, where $\mathbf{x}$ denotes the input image and $y$ denotes the associated class label. Ideally, we want to train a model by minimizing the empirical risk on the union of datasets of all the clients, $\mathcal{D}=\cup_{n\in N}\mathcal{D}_{n}$. However, due to the distributed nature of the datasets in the FL setting, we can only train a model on the local data of the client, ie, conduct local training, as:
$\min_{\theta}\mathcal{L}(\theta)=\sum^{|D_{n}|}_{i=1}\mathcal{L}_{\text{ce}}(f_{\theta}(\mathbf{x}^{(i)}),y^{(i)}),$ (1)
where $\theta$ are the parameters of the $n$-th client model $f_{\theta}$ and $\mathcal{L}_{\text{ce}}$ is a supervised cross-entropy loss.
To aggregate knowledge (or global aggregation), the FL setting involves clients communicating the weights of the trained model $f_{\theta}$ to the server through multiple rounds of communication. In each aggregation round $t$, the server aggregates client weights via a weighted average, or more concretely using the FedAvg [fedavg] as:
$\theta_{t}=\sum^{N}_{n=1}\frac{|D_{n}|}{|D|}\theta^{(n)}_{t}$, (2)
@@ 3.1 Federated Learning
3.2 Federated fine-tuning of CLIP
Textual prompt tuning (TPT), ie, CoOp [coop], is a commonly used fine-tuning approach to improve the classification accuracy of CLIP on downstream tasks. For fine-tuning, TPT learns a set of prompts while keeping the weights of the pretrained CLIP image encoder $\mathcal{E}_{v}$ and text encoder $\mathcal{E}_{t}$ frozen. In detail, TPT replaces handcrafted prompts (eg, “a photo of a [CLS]”) at the input of the text encoder $\mathcal{E}_{t}$ with a set of lightweight learnable vectors $\mathcal{T}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M}\}$, where $M$ denotes the length of the tokens. Text features are computed with the prompt $\mathbf{t}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M},[\text{CLS}]\}$, and the image features are computed from the image $\mathbf{x}$, using the respective encoders. The prompts $\mathcal{T}$ are learned using the standard cross-entropy loss, as in Eq. (1).
Since the prompts $\mathcal{T}$ in TPT are lightweight and incur low communication overhead, it has become a dominant approach for fine-tuning CLIP in the FL setting [promptfl, diprompt, fedclip]. Each client trains prompts $\mathcal{T}$ on the local client data and communicates them with the server. The server aggregates the prompts $\mathcal{T}$ from different clients using Eq. (2), and sends back the aggregated prompts back to the clients. TPT-based FL methods differ mostly in how the prompts are designed, trained locally, and aggregated on the server.
While TPT is not the only approach to fine-tune CLIP in the FL setting, it remains the most dominant approach. Works in related areas have explored using Adapters [gao2024clip] for fine-tuning CLIP in the FL setting [fedclip], and LoRA [hu2022lora] for fine-tuning LLMs in a multi-task [bohdal2025efficient] and FL setting [yang2025federated].
3.3 Model calibration
Besides accuracy, we also want a model to be well-calibrated. That is, we want the predicted class probabilities of the model to be an accurate representation of the model’s accuracy [guo2017calibration]. Formally, a model is said to be well-calibrated (or confidence-calibrated), if for a confidence of $c$, the model is correct $c$ proportions of the time:
@@ 3.3 Model calibration
$\text{ECE}=\sum^{G}_{g=1}\frac{\mid b_{g}\mid}{n}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid,$
where $\text{acc}(\cdot)$ and $\text{conf}(\cdot)$ denotes the average accuracy and average confidence of the model for the samples falling in the bin $b_{g}$, respectively. A well-calibrated model will have a low ECE.
In the context of a centralized setting, it has been shown in the literature that although zero-shot CLIP (ZS-CLIP) is well-calibrated [minderer2021revisiting] for the few-shot image classification setting, fine-tuning can make a model miscalibrated [wang2024open, dor]. In particular, during CLIP fine-tuning, a regularization term (eg, regularizing text encoder divergence [dor]) is added by the in-training calibration methods to prevent the model from being miscalibrated. Although pertinent, no prior work has studied the impact of fine-tuning CLIP on the model’s calibration in a FL setting.
@@ 4.1 Experimental setup
Benchmarks and settings. For evaluation, we use 10 benchmarks: CIFAR-10, CIFAR-100 [cifar], OfficeHome [officehome], PACS [pacs], VLCS [vlcs], Food101 [food], DTD [dtd], Caltech101 [caltech], Flowers102 [oxfordflowers], and OxfordPets [oxfordpets]. We follow the non-IID FL setting, where the clients have training data of variable number of classes and the classes can be shared among the clients, ie, overlapping classes. Following previous work [cao2023knowledge, shamsian2021personalized], we simulate data heterogeneity by partitioning the data among the clients by setting the parameter $\alpha=0.5$ of a symmetric Dirichlet distribution.
@@
We group the benchmarks in a manner to simulate different characteristics of a heterogeneous setup: (i) in-distribution setting where only a single dataset (eg, CIFAR-100) is distributed among the clients, (ii) domain generalization setting where different domains of the same benchmark are distributed among the clients (eg, the four domains of OfficeHome), and (iii) a base-to-new generalization setting where some of the classes are held out during training to measure generalization on unseen classes. Details of different settings have been reported in the Appendix.
Baselines. In this study, we consider a range of TPT methods proposed for the task of image classification with CLIP. In particular, we consider the following methods:
[-] CoOp [coop], the seminal paper that proposed TPT for image classification in the centralized setting. In addition, PromptFL [promptfl] that extends CoOp to the FL setting.
[-] KgCoOp [kgcoop] that adds a regularization term to CoOp optimization, and its FL variant FedKgCoOp.
[-] MaPLe [maple] that extends TPT also to the vision encoder, and its FL variant FedMaPLe.
[-] State-of-the-art TPT methods that have been specifically designed to operate in the personalized FL setting: FedPGP [fedpgp], FedOTP [fedotp] and FedPHA [fedpha].
For improving calibration off-the-shelf, we have used three widely used approaches: temperature scaling [guo2017calibration], which is a post-hoc calibration technique; and two in-training calibration techniques, DCA [dca] and MDCA [mdca].
@@
Implementation Details. In our experiments, we employ the frozen ViT-B/16 variant of CLIP backbone. We consider the rank of LoRA metrices to $2$ and a dropout rate of $0.25$ to regularize the input of the LoRA modules. We keep $\alpha=0.5$ of a Dirichlet distribution, communication rounds 50 with local epoch 1 with participation rate of 100% for all datasets, except CIFAR-10/100 with 10% participation rate. We choose 16 shots per class across over 10 clients for base-to-new generalization task, two clients per domain in domain generalization and 100 clients per dataset on in-distribution setting. We inherit the optimizer setup, using SGD [sgd] with a learning rate of $1\times 10^{-3}$, warm-up constant of $1\times 10^{-5}$ and a batch size of $32$.
@@
Evaluation metrics. To compare methods, we use the following 4 metrics: classification accuracy (Acc $\uparrow$), Expected Calibration Error (ECE $\downarrow$) [guo2017calibration], Maximum Calibration Error (MCE $\downarrow$) [guo2017calibration], and Adaptive Calibration Error (ACE $\downarrow$) [nixon2019measuring]. We report performances in the personalized FL setting [fedotp, fedpgp, fedpha], where the metrics are computed at each client and then averaged across clients.
4.2 Calibration error in federated learning
@@
Table 2: Comparison of offline CLIP fine-tuning with federated CLIP fine-tuning. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100.
Method CIFAR-10 CIFAR-100
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
ZS-CLIP [radford2019language] 89.91 10.16 2.62 8.54 66.84 15.06 3.84 14.98
Offline setting (without calibration)
CoOp [coop] 93.99 1.00 0.56 0.89 77.86 1.93 0.57 1.77
KgCoOp [kgcoop] 92.17 2.70 0.61 2.70 74.73 7.83 1.13 7.83
MaPLe [maple] 97.08 9.05 1.45 1.81 85.01 5.42 2.04 4.72
Offline setting (with calibration: DCA [dca])
CoOp 93.89 1.20 0.62 0.87 77.85 1.83 0.73 1.71
KgCoOp 92.37 2.21 0.46 2.21 74.82 5.30 0.80 5.30
MaPLe 96.96 10.93 1.56 2.10 84.99 5.91 2.11 5.17
FL setting (without calibration)
FedCoOp [promptfl] 92.42 4.94 1.64 3.94 72.43 9.27 2.43 8.66
FedKgCoOp 90.48 6.53 1.88 5.79 70.68 11.20 2.95 11.05
FedMaPLe 96.08 3.49 1.28 2.18 81.73 7.53 2.20 6.93

### v2

@@ 1 Introduction
Reliability is a critical consideration when deploying deep learning (DL) models in production scenarios. We aim to develop DL models that are not only highly accurate but also reliable. Reliability has many facets, including robustness to covariate shift [csurka2017domain], model calibration [guo2017calibration], and generalization to out-of-distribution data [yang2024generalized]. Of particular interest, model calibration is crucial because it ensures that a model’s predicted probabilities accurately reflect the true likelihood of correctness, enabling reliable decision-making [pavlovic2025understanding]. As powerful DL models (e.g., CLIP [radford2019language]) are increasingly being deployed in high-stake applications, such as healthcare and autonomous systems, ensuring proper calibration is crucial.
Figure 1: Comparison of reliability diagrams and calibration errors. Expected Calibration Error (ECE $\downarrow$) [ece] is reported on the OxfordPets [oxfordpets] dataset on a centralized training or offline (left) and a non-IID personalized federal learning (FL) setting (right).
Table 1: Overview of works focusing on model calibration. While previous works have studied calibration of CLIP, no work has considered how fine-tuning CLIP in the FL setting affects calibration. Our work explores this direction.
Method CLIP Calibration Fine-tune FL
@@
In this work, we make an attempt to answer the question: how much calibrated is CLIP when fine-tuned in a Federated Learning setup? Towards that goal, we first analyze if the trend of miscalibration, observed in the offline CLIP fine-tuning setting [dor, wang2024open], also holds true in the FL setting. As shown in Fig. 1, CoOp [coop], a textual-prompt tuning approach, when fine-tuned in the FL setup (PromptFL [promptfl]) is even more miscalibrated. Imposing regularization through in-training calibration (e.g., with DCA [dca]) to both the offline and FL setting alleviates miscalibration by some extent. This observation conforms with the findings of [nucfl], where a similar behaviour is shown in the FL setup.
However, as shown in Tab. 1, the work in [nucfl] only experimented with the classical ResNet family of models trained from scratch and did not explore a multimodal setting. Differently, in our study we are specifically interested in analyzing the miscalibration of fine-tuned CLIP models in the FL setup, which exhibits substantially different dynamics from standard Convnets [minderer2021revisiting]. In Sec. 4, we dive deep into this issue, by investigating existing methods for improving calibration, and our empirical analyses show that improving calibration of CLIP-based methods in the FL setup is more challenging than the offline setting.
Core findings of our work Fine-tuning CLIP in the FL setting is more prone to miscalibration than in the centralized setting Fine-tuning strategy affect a lot the calibration, with LoRA being more effective than other PEFT methods Explicitly optimizing for calibration rarely produces systematic improvements, and often can be detrimental
@@
As an alternative, we strive to find a simple approach that not only preserves the accuracy of the model, but also ensures a well-calibrated model without introducing additional regularization. To this end, we deviate from the common approach of textual prompt-tuning in FL [promptfl, diprompt, fedotp], and adopt low rank adaptation (LoRA [lora]) as a means to fine-tune CLIP. We call our approach $\text{FL}^{2}\text{oRA}$ (Federated Learning with LoRA). As shown in Fig. 1 (right), $\text{FL}^{2}\text{oRA}$ alone is better calibrated than PromptFL + DCA, without the need for any explicit regularization on the network. With extensive experiments on three FL settings that encompass generalization to new classes and domains and their combination, we show that fine-tuning CLIP with LoRA alone is highly effective in preserving model calibration in the FL setting, and adding extra regularization to improve calibration brings only a marginal improvement. Note that the goal of this work is not to introduce a novel architecture for FL, but to fill the gap in the literature by empirically analyzing calibration in federated CLIP fine-tuning and proposing a simple, while effective approach to preserve calibration.
@@
Contributions. In summary: (i) We are the first to empirically study the impact of fine-tuning CLIP on the model’s calibration in the FL setup. (ii) We present $\text{FL}^{2}\text{oRA}$, a simple and effective approach, that departs from the dominant paradigm of fine-tuning CLIP with textual prompt-tuning. (iii) We conduct extensive experiments across diverse FL scenarios and show that our LoRA-based federated fine-tuning consistently provides a strong baseline, potentially removing the need for additional regularization for calibration.
@@ 2 Related Work
Federated learning enables collaborative model training across decentralized clients without sharing local data. A core challenge is client heterogeneity, e.g., non-IID data distributions, which can significantly degrade performance. Prior work primarily focuses on designing robust aggregation and optimization strategies to address this issue. Representative methods include FedAvg [fedavg], which averages client updates, FedProx [fedprox], which introduces a proximal term to mitigate client drift and FedNova [fednova], which normalizes updates to improve convergence under heterogeneous objectives. Broadly, these approaches aim to improve the accuracy and convergence of the aggregated global model. In contrast, we investigate a relatively underexplored aspect of FL: its impact on model calibration in Vision Language Models (VLMs).
@@
Federated fine-tuning of CLIP. When applied to large VLMs such as CLIP [radford2019language], FL introduces substantial computational and communication overhead due to the need to fine-tune millions of parameters. To mitigate this, prompt tuning [coop, kgcoop, vpt, maple] has emerged as an efficient alternative for adapting CLIP. Recent works extend prompt tuning to the federated setting [bai2024diprompt, promptfl, qiu2024federated, sun2024towards, fedotp, fedtpg, fedpgp, fedpha]. For example, PromptFL [promptfl] regularizes client-specific prompts, while FedOTP [fedotp] and FedTPG [fedtpg] improve communication efficiency and personalization via orthogonal or task-conditioned prompt generation. FedCLIP [fedclip] decouples global and local adaptation of the text and image encoders, and methods such as FedTPG [fedtpg] and FedMVP [fedmvp] dynamically generate prompts using auxiliary information (e.g., class names or visual attributes).
@@
In contrast, we adopt a Parameter-Efficient Fine-Tuning (PEFT) approach, i.e., LoRA [biderman2024lora], to adapt CLIP in the federated setting. While FL with LoRA has been studied for LLMs [yang2025federated], prior work focuses mainly on accuracy. To our knowledge, we are the first to study calibration in CLIP-based federated image classification.
@@
Model calibration [pavlovic2025understanding] measures how well predicted confidence scores reflect true correctness likelihood. Calibration methods are generally divided into post-hoc and training-time approaches. Post-hoc methods, such as temperature scaling [guo2017calibration] and Dirichlet calibration [kull2017beta], adjust output probabilities after training without modifying model parameters, offering simplicity but limited flexibility. In contrast, training-time methods, including label smoothing [muller2019does], focal loss [lin2017focal], DCA [dca], and MDCA [hebbalaguppe2022stitch], modify the objective to encourage well-calibrated predictions, often at the cost of additional hyperparameter tuning.
@@
For VLMs, zero-shot CLIP is known to be well-calibrated [minderer2021revisiting], but fine-tuning can degrade calibration, motivating approaches such as DOR [dor], DAC [wang2024open], and CalShift [khan2025confidence]. Most prior studies focus on offline settings. In FL, only a few works [peng2024fedcal, nucfl] address calibration, primarily with fully fine-tuned CNNs. To our knowledge, calibration in CLIP-based federated learning remains unexplored.
@@ 3 Preliminaries
In this section, we discuss the preliminaries related to federated learning, federated fine-tuning of CLIP and model calibration.
Federated learning. For the task of image classification, the FL scheme consists of training data distributed among $N$ clients and each client $n$ possess a dataset $\mathcal{D}_{n}=\{\mathbf{x}^{(i)},y^{(i)}\}^{\mid\mathcal{D}_{n}\mid}_{i=1}$, where $\mathbf{x}^{(i)}$ denotes the $i$-th input image and $y^{(i)}$ denotes the class label. Ideally, we want to train a model by minimizing the empirical risk on the union of datasets of all the clients, $\mathcal{D}=\cup_{n\in N}\mathcal{D}_{n}$. However, due to the distributed nature of the datasets in the FL setting, we can only train a model $f_{\theta}$ on the local data of the client, i.e., conduct local training, as:
$\min_{\theta}\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{n}|}\sum^{|\mathcal{D}_{n}|}_{i=1}\mathcal{L}_{\text{ce}}(f_{\theta}(\mathbf{x}^{(i)}),y^{(i)}),$ (1)
where $\theta$ are the weights of a client model and $\mathcal{L}_{\text{ce}}$ is cross-entropy loss.
To aggregate knowledge (or global aggregation), the FL setting involves clients communicating the weights of the trained model $f_{\theta}$ to the server through multiple rounds of communication. In each round $t$, the server aggregates client weights via a weighted average, or, more concretely, using FedAvg [fedavg] as:
$\theta_{t}=\sum^{N}_{n=1}\frac{|\mathcal{D}_{n}|}{|\mathcal{D}|}\theta^{(n)}_{t}$, (2)
@@
Federated fine-tuning of CLIP. Textual prompt tuning (TPT), i.e., CoOp [coop], is a commonly used fine-tuning approach to improve the classification accuracy of CLIP on downstream tasks. For fine-tuning, TPT learns a set of prompts while keeping the weights of the pretrained CLIP image encoder $\mathcal{E}_{v}$ and text encoder $\mathcal{E}_{t}$ frozen. In detail, TPT replaces handcrafted prompts (e.g., “a photo of a [CLS]”) at the input of the text encoder $\mathcal{E}_{t}$ with a set of lightweight learnable vectors (or tokens) $\mathcal{T}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M}\}$, where $M$ denotes the length of the tokens. Text features are computed with the prompt $\mathbf{t}=\{\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{M},[\text{CLS}]\}$, and the image features are computed from the image $\mathbf{x}$, using the respective encoders. The prompts $\mathcal{T}$ are learned using cross-entropy loss, as in Eq. (1).
Since the prompts $\mathcal{T}$ in TPT are lightweight and incur low communication overhead, it has become a dominant approach for fine-tuning CLIP in the FL setting [promptfl, diprompt, fedclip]. Each client trains prompts $\mathcal{T}$ on the local client data and communicates them with the server. The server aggregates the prompts $\mathcal{T}$ from different clients using Eq. (2), and sends back the aggregated prompts to the clients. TPT-based FL methods differ mostly in how the prompts are designed, trained locally, and aggregated on the server.
While TPT is not the only method for fine-tuning CLIP in FL, it remains the most widely used. Related work has explored Adapters for FL of CLIP [fedclip] and LoRA for fine-tuning LLMs in multi-task and federated settings [bohdal2025efficient, yang2025federated].
Model calibration. In addition to higher accuracy, we also want a model to be well-calibrated. That is, we want the predicted class probabilities of the model to be an accurate representation of the model’s accuracy [guo2017calibration]. Formally, a model is said to be well-calibrated (or confidence-calibrated), if for a confidence of $c$, the model is correct $c$ proportions of the time:
@@
$\text{ECE}=\sum^{G}_{g=1}\frac{\mid b_{g}\mid}{|\mathcal{D}|}\mid\text{acc}(b_{g})-\text{conf}(b_{g})\mid,$
where $\text{acc}(b_{g})$ and $\text{conf}(g_{g})$ denotes the average accuracy and average confidence of the model for the samples falling in the bin $b_{g}$, respectively. A well-calibrated model will have a low ECE.
In centralized (offline) settings, prior work showed that while zero-shot CLIP is well-calibrated in few-shot classification [minderer2021revisiting], fine-tuning can degrade calibration [wang2024open, dor]. To mitigate this, methods introduce regularization (e.g., constraining text encoder divergence [dor]) during fine-tuning. However, the impact of fine-tuning CLIP on calibration in FL remains unexplored.
@@ 4.1 Experimental setup
Benchmarks and settings. For extensive evaluation, we use a collection of 10 benchmarks: CIFAR-10, CIFAR-100 [cifar], OfficeHome [officehome], PACS [pacs], VLCS [vlcs], Food101 [food], DTD [dtd], Caltech101 [caltech], Flowers102 [oxfordflowers], and OxfordPets [oxfordpets]. We follow the non-IID FL setting, where the clients have training data of variable number of classes and the classes can be shared among the clients, i.e. overlapping classes. Following previous work [cao2023knowledge, shamsian2021personalized], we simulate data heterogeneity by partitioning the data among the clients by setting the parameter $\beta=0.5$ of a symmetric Dirichlet distribution.
@@
We organize the benchmarks to simulate heterogeneous FL scenarios: (i) an in-distribution setting where a single dataset (e.g., CIFAR-100) is split across clients, (ii) a domain generalization setting where different domains (e.g., OfficeHome) are assigned to different clients, and (iii) a base-to-new generalization setting where some classes are held out during training. Additional details are provided in the Appendix.
Baselines. We evaluate several TPT methods for CLIP-based image classification: (i) PromptFL [promptfl], which extends CoOp [coop] to FL; (ii) KgCoOp [kgcoop], which adds a regularization term to CoOp, along with its federated variant, FedKgCoOp; (iii) state-of-the-art TPT methods for personalized FL: FedPGP [fedpgp], FedOTP [fedotp], and FedPHA [fedpha]. For off-the-shelf calibration, we use temperature scaling [guo2017calibration] (post-hoc method) and two training-based methods: DCA [dca] and MDCA [mdca].
@@
Implementation Details. We use a frozen CLIP ViT-B/16 backbone. We set the Dirichlet parameter to $\beta=0.5$, run 50 communication rounds with 1 local epoch, and use 100% participation rate (10% for CIFAR-10/100). For base-to-new generalization, we use 16 shots per class across 10 clients; for domain generalization, two clients per domain; and for in-distribution settings, 100 clients per dataset. We follow the standard optimizer setup with SGD [sgd], a learning rate of $1\times 10^{-3}$, warm-up $1\times 10^{-5}$, and batch size 32.
@@
Evaluation metrics. To compare methods, we compute: classification accuracy (Acc $\uparrow$), Expected Calibration Error (ECE $\downarrow$) [guo2017calibration], Maximum Calibration Error (MCE $\downarrow$) [guo2017calibration], and Adaptive Calibration Error (ACE $\downarrow$) [nixon2019measuring]. We report performances in the personalized FL setting [fedotp, fedpgp, fedpha], where the metrics are computed at each client and then averaged across clients.
@@
Table 2: Comparison of offline CLIP fine-tuning with federated CLIP fine-tuning. Results have been reported on the in-distribution setting using datasets: CIFAR-10 and CIFAR-100. Best number is in bold.
Method Offline setting (w/o calib) Offline setting (w/ calib: DCA [dca]) FL setting (w/o calib)
Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$ Acc. $\uparrow$ ECE $\downarrow$ MCE $\downarrow$ ACE $\downarrow$
CIFAR-10
CoOp [coop] 93.99 1.00 0.56 0.89 93.89 1.20 0.62 0.87 92.42 4.94 1.64 3.94
KgCoOp [kgcoop] 92.17 2.70 0.61 2.70 92.37 2.21 0.46 2.21 90.48 6.53 1.88 5.79
VPT [vpt] 96.93 1.23 0.64 0.90 96.98 1.45 0.75 0.94 95.29 3.53 1.22 2.52
CIFAR-100
CoOp [coop] 77.86 1.93 0.57 1.77 77.85 1.83 0.73 1.71 72.43 9.27 2.43 8.66
KgCoOp [kgcoop] 74.73 7.83 1.13 7.83 74.82 5.30 0.80 5.30 70.68 11.20 2.95 11.05
VPT [vpt] 80.86 1.45 0.41 1.35 80.82 1.47 0.39 1.46 77.83 7.94 2.39 8.00
4.2 How does Federated Learning affect calibration?